# TVR: Automotive System Requirement Traceability Validation and Recovery Through Retrieval-Augmented Generation

### v1

@@ Abstract.
In automotive software development, as well as other domains, traceability between stakeholder requirements and system requirements is crucial to ensure consistency, correctness, and regulatory compliance. However, erroneous or missing traceability relationships often arise due to improper propagation of requirement changes or human errors in requirement mapping, leading to inconsistencies and increased maintenance costs. Existing approaches do not address traceability between stakeholder and system requirements, rely on open-source data—as opposed to automotive (or any industry) data—and do not address the validation of manual links established by engineers. Additionally, automotive requirements often exhibit variations in the way they are expressed, posing challenges for supervised models requiring training. The recent advancements in large language models (LLMs) provide new opportunities to address these challenges. In this paper, we introduce TVR, a requirement Traceability Validation and Recovery approach primarily targeting automotive systems, leveraging LLMs enhanced with retrieval-augmented generation (RAG). TVR is designed to validate existing traceability links and recover missing ones with high accuracy. We empirically evaluate TVR on automotive requirements, achieving 98.87% accuracy in traceability validation and 85.50% correctness in traceability recovery. Additionally, TVR demonstrates strong robustness, achieving 97.13% in accuracy when handling unseen requirements variations. The results highlight the practical effectiveness of RAG-based LLM approaches in industrial settings, offering a promising solution for improving requirements traceability in complex automotive systems.
copyright: nonecopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; xxx; xxxisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Maintaining accurate and high-quality traceability links between stakeholder and system requirements is indeed crucial for multiple reasons: 1) Enhancing system consistency; ensuring that requirements are correctly implemented minimizes functional errors and inconsistencies (Wiegers and Beatty, 2013; Pargaonkar, 2023). 2) Preventing error propagation – Detecting incorrect requirement mappings early reduces maintenance costs and avoids large-scale rework (Tufail et al., 2017). 3) Ensuring regulatory compliance, thus complying with ASPICE standards to help improve the safety, reliability, and auditability of automotive software (Qusef et al., 2011).
@@
However, significant challenges have been reported by industry engineers regarding traceability between requirements, for example in the context of Diagnostic Trouble Code (DTC) management, which refer to standardized vehicle error codes indicating malfunctions detected by the diagnostics system (Marscholik and Subke, 2009; Pirasteh et al., 2019). According to system engineers, the primary causes of these challenges are related to resource and time constraints, and include: 1) Improper propagation of requirement changes: When stakeholder requirements are modified (e.g., new requirements are added or existing ones are updated), the corresponding system requirements may not be updated accordingly, leading to outdated or invalid traceability links. If a requirement is deleted or merged, the original traceability link may persist, resulting in incorrect mappings that could cause functional inconsistencies or safety risks. 2) Errors in requirement mapping: Due to human error, system requirements may be incorrectly mapped to unrelated stakeholder requirements, or traceability links are missing.
The goal of this paper is to provide automated, effective means to address the challenges above and thus identify invalid and missing traceability links between DTC requirements. Because requirements are not always expressed consistently within or across projects, we rely on large language models (LLMs), hoping they are robust to such variations.
We are motivated in our choice by the fact that LLMs have demonstrated significant advancements in natural language processing (NLP) (Gao et al., 2024; Karanikolas et al., 2023; Hu and Lu, 2024; Tang et al., 2023) and software engineering tasks (Gao et al., 2024; Hu and Lu, 2024; Karanikolas et al., 2023), especially code generation (Du et al., 2024; Mu et al., 2024; Gao et al., 2023a; Chen et al., 2024; Zhang et al., 2024; Tang et al., 2023), and code summarization (Nam et al., 2024; Ahmed et al., 2024; Geng et al., 2024). Further, LLMs have also shown to be effective for traceability link recovery (Fuchß et al., 2025; Rodriguez et al., 2023).
@@
To address traceability issues and meet the practical needs of our industrial partner, we propose a Retrieval-Augmented Generation (RAG) enhanced approach: Traceability Validation and Recovery (TVR). We investigate and leverage many state-of-the-art (SOTA) LLMs—with Claude Sonnet 3.5 (Anthropic, 2024) yielding the best results—to verify whether a traceability link is valid and recover missing links between requirements. We evaluated TVR on 2,312 DTC requirement pairs from an automotive system, achieving an overall accuracy of 98.87%. To investigate its robustness against variations in the way requirements are written (as detailed in Section 2), we evaluated TVR on new requirement variations. The results show that TVR still maintains 97.13% in accuracy, demonstrating strong robustness. Additionally, to address the challenge of missing links, we applied TVR to traceability link recovery, successfully identifying 502 missing traceability links with an accuracy of 85.50%.
@@
  • Proposing TVR, a RAG-enhanced LLM approach for validating traceability links between automotive requirements.
  • Evaluating the robustness of TVR on unseen requirement variations.
  • Applying TVR to recover traceability links.
  • Investigating 13 SOTA LLMs across five prompt engineering strategies, as well as the impact of different similarity measures and the number of examples on RAG performance.
@@ 2.1.2. DTC Stakeholder Requirements
Figure 1. Fictitious Variations of Stakeholder Requirements
@@ 2.1.3. DTC System Requirements
Figure 2. Example of Sanitized DTC System Requirement
@@ 2.2. Requirements Traceability
To address these two challenges, this paper proposes TVR, a method for verifying requirements traceability for automotive systems. TVR leverages the exceptional natural language understanding capabilities of LLMs. Instead of relying on the general concept of traceability links (Guo et al., 2024), we ask LLMs to verify whether our system requirement covers the message or signal specified in the stakeholder requirement. We guide the LLM to focus specifically on the message or signal, rather than other parts of the requirement, to avoid confusion.
@@ 3. TVR
The overall workflow of TVR, illustrated in Figure 3, consists of two key components: a retriever and a generator. The retriever identifies the most similar examples from $D_{\text{train}}$ based on their similarity to $x_{\text{test}}$ and provides them as input to the generator. The generator then utilizes these retrieved examples to assess the validity of the traceability link between $stake_{\text{test}}$ and $sysReq_{\text{test}}$.
@@ 3. RQ3:
    How robust is TVR for requirements traceability validation in the presence of unseen requirements variations?
    One of our main motivations in relying on LLMs is that we expect them to be more robust to such variations. Due to the numerous variations that are typically present in stakeholder requirements, this RQ primarily evaluates the performance of TVR across unseen variations in stakeholder requirements.
@@ 4.2.1. LLMs
In this study, due to our industry partner’s data privacy concerns, we were unable to use LLMs such as the GPT series (OpenAI, 2023a, b). Instead, we were provided with 13 SOTA LLMs available through the Amazon Bedrock service, which include Claude, Llama, Mistral, and Titan. Extensive experiments were conducted using these LLMs.
@@ 4.2.2. Prompt Engineering
  • Zero-shot Prompting (Brown et al., 2020). Zero-shot prompting directly instructs the model to complete a task without providing demonstrations or additional context.
  • Chain-of-Thought Prompting (CoT) (Wei et al., 2022). CoT enhances the complex reasoning capabilities of LLMs by incorporating intermediate reasoning steps.
@@
  • Retrieval-Augmented Generation (RAG) (Lewis et al., 2020). RAG enhances LLMs’ capability of solving knowledge-intensive tasks by integrating information retrieval with text generation. It retrieves relevant external knowledge sources to provide additional context, enabling models to generate more accurate and up-to-date responses without requiring full retraining (Gao et al., 2023b; Riedel et al., 2020).
Note that in our experiments, each prompting strategy builds upon the previous one. The CoT reasoning extends Zero-Shot, while few-shot prompting enhances CoT by incorporating examples in the prompt. The self-consistency strategy further refines few-Shot prompting by applying a majority voting method. RAG improves few-shot prompting by selecting more relevant examples.
@@ 4.3.1. Data Collection
To evaluate the performance of LLMs in autonomous requirements traceability validation and recovery, we collected a dataset of DTC requirements from our industry partner. We began by gathering high-quality DTC system requirements, including the “Lost Communication” and “Implausible Data” DTC system requirements, which are among the most common and critical types of DTC requirements. Subsequently, we identified the stakeholder requirements linked to these system requirements, yielding a total of 1,320 stakeholder requirements.
@@ 4.3.2. Data Cleaning
To check the completeness of stakeholder requirements, we defined regular expressions to validate their structure 777while we acknowledge that more advanced approaches for completeness checking exist, this is not the primary focus of our study. Specifically, each atomic requirement must include a trigger condition, an input message, and a mitigation action. We filtered out both stakeholder and system requirements that did not meet the above two constraints. In practice, such requirements should be flagged as issues requiring correction and must be addressed before performing requirement traceability validation.
@@ 4.4.1. Traceability Validation
To assess the performance of TVR in autonomous requirements traceability validation, we employ standard metrics, including accuracy, precision, recall, and F1 score. $TP$ is the number of traceability links that are correctly identified as valid. $FN$ is the number of traceability links that are incorrectly identified as invalid. $FP$ is the number of traceability links that are incorrectly identified as valid. $TN$ is the number of traceability links that are correctly identified as invalid.
Accuracy is the proportion of correctly identified traceability links (both valid and invalid) out of all traceability links. Precision is the proportion of correctly identified valid links ($TP$) out of all links identified as valid by the model ($TP$ + $FP$). It measures how reliable the model’s predictions of valid links are. Recall is the proportion of correctly identified valid links ($TP$) out of all actual valid links in the ground truth ($TP$ + $FN$). It measures the model’s ability to find all the valid links. F1 score is the harmonic mean of precision and recall. It balances the trade-off between precision and recall, providing a single measure of the model’s performance.

### v2

# TVR: Automotive System Requirements Traceability Validation and Recovery Through Retrieval-Augmented Generation

@@ Abstract.
In automotive software development, as well as other domains, traceability between stakeholder requirements and system requirements is crucial to ensure consistency, correctness, and regulatory compliance. However, erroneous or missing traceability relationships often arise due to improper propagation of requirement changes or human errors in requirement mapping, leading to inconsistencies and increased maintenance costs. Existing approaches do not address traceability between stakeholder and system requirements, rely on open-source data—as opposed to automotive (or any industry) data—and do not address the validation of manual links established by engineers. Additionally, automotive requirements often exhibit variations in the way they are expressed, posing challenges for supervised models requiring training. The recent advancements in large language models (LLMs) provide new opportunities to address these challenges. In this paper, we introduce TVR, a requirement Traceability Validation and Recovery approach primarily targeting automotive systems, leveraging LLMs enhanced with retrieval-augmented generation (RAG). TVR is designed to validate existing traceability links and recover missing ones with high accuracy. We empirically evaluate TVR on automotive requirements, achieving 98.87% accuracy in traceability validation and 85.50% correctness in traceability recovery. Additionally, TVR demonstrates strong robustness, achieving 97.13% in accuracy when handling unseen requirement variations. The results highlight the practical effectiveness of RAG-based LLM approaches in industrial settings, offering a promising solution for improving requirements traceability in complex automotive systems.
Requirements Engineering, Traceability Validation and Recovery, Large Language Models, Automotive System
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Requirements analysisccs: Software and its engineering Traceability
@@ 1. Introduction
Maintaining accurate and high-quality traceability links between stakeholder and system requirements is indeed crucial for multiple reasons: 1) Improving system consistency—Consistent traceability ensures that system requirements accurately reflect stakeholder needs, thus minimizing functional errors and inconsistencies (Wiegers and Beatty, 2013; Pargaonkar, 2023). 2) Preventing error propagation—Detecting incorrect requirement mappings early reduces maintenance costs overhead and avoids large-scale rework (Tufail et al., 2017). 3) Ensuring regulatory compliance, thus complying with ASPICE and functional-safety standards—including ISO 26262 (Siegl et al., 2010)to help improve the safety, reliability, and auditability of automotive software (Qusef et al., 2011).
@@
However, significant challenges have been reported by industry engineers regarding traceability between requirements. One specific and particularly critical type of requirements, which we focus on here, are related to Diagnostic Trouble Code (DTC) management. DTCs refer to standardized vehicle error codes indicating malfunctions detected by the diagnostics system (Marscholik and Subke, 2009; Pirasteh et al., 2019). While automotive diagnostic tools—with the market valued at USD 38.45 billion in 2023 and projected to reach USD 56.07 billion by 2031—play a crucial role in early issue detection, maintaining correct traceability between DTC requirements remains a largely manual and labor-intensive task. This limits the overall efficiency of the diagnostic process due to the time and efforts required to manually verify the consistency between requirements and trace the system faults to requirements. According to system engineers, these challenges, primarily due to resource and time constraints, include: 1) Improper propagation of requirement changes: When stakeholder requirements are modified (e.g., new requirements are added or existing ones are updated), the corresponding system requirements may not be updated accordingly, leading to outdated or invalid traceability links. If a requirement is deleted or merged, the original traceability link may persist, resulting in incorrect mappings that could cause functional inconsistencies or safety risks—potentially causing irrecoverable losses in safety-critical domains. For example, in 1998, NASA’s \$327 M Mars Climate Orbiter  [^2] https://en.wikipedia.org/wiki/Mars_Climate_Orbiter failed because mission directives mandated SI units, but supporting software output imperial units. Without end-to-end traceability, the mismatch went undetected, causing trajectory errors and mission loss. This underscores the need for automated, robust requirement-consistency validation in industrial systems. 2) Errors in requirement mapping: Due to human error, system requirements may be incorrectly mapped to unrelated stakeholder requirements, or traceability links are missing. According to our industry partner, they spend several weeks each year on manually verifying requirements traceability and consistency, as ensuring compliance with stakeholder requirements is always a top priority. Automating this process could significantly reduce their manual effort and associated labor costs.
Unlike similarity-based approaches (Gao et al., 2022; Hey et al., 2021; Panichella et al., 2013; Kuang et al., 2015; Antoniol et al., 2002; Marcus and Maletic, 2003; Keim et al., 2024, 2023; Cleland-Huang et al., 2005) that reconstruct links between requirements from scratch, industrial settings typically rely on preexisting traceability links—which, due to labeling errors or version changes, can incorrectly tie together unrelated or diverged requirements. Hence, it is essential to validate these links to ensure the connected requirements remain semantically and content-wise consistent.
The goal of this paper is to provide automated, effective ways to address the challenges above and thus identify invalid and missing traceability links between DTC requirements. Because requirements are not always expressed consistently within or across projects, we leverage large language models (LLMs), for their demonstrated robustness to linguistic variation. LLMs have recently achieved significant advancements in natural language processing (NLP) (Gao et al., 2024; Karanikolas et al., 2023; Hu and Lu, 2024; Tang et al., 2023) and software engineering tasks (Gao et al., 2024; Hu and Lu, 2024; Karanikolas et al., 2023), especially code generation (Du et al., 2024; Mu et al., 2024; Gao et al., 2023a; Chen et al., 2024; Zhang et al., 2024; Tang et al., 2023), and code summarization (Nam et al., 2024; Ahmed et al., 2024; Geng et al., 2024). Although prompt‐based LLM approaches have been recently applied to traceability recovery (Fuchß et al., 2025; Rodriguez et al., 2023), they either rely solely on simple zero-shot prompts or RAG-based retrieval followed by generic queries (e.g., “Is there a traceability link between”), which are inadequate for precisely verifying the links between DTC system and stakeholder requirements that tend to slightly differ in the character strings of messages and signals.
@@
To address traceability challenges in DTC requirements described above, a Retrieval-Augmented Generation (RAG) enhanced approach that leverages RAG to retrieve similar requirement pairs to be included in the prompts. We investigate and leverage many state-of-the-art (SOTA) LLMs—with Claude Sonnet 3.5 (Anthropic, 2024) yielding the best results—to verify whether a traceability link is valid and recover missing links between requirements. We evaluated TVR on 2,132 DTC requirement pairs from an automotive system, achieving an overall accuracy of 98.87%. To investigate its robustness against variations in the way requirements are written (as detailed in Section 2), we evaluated TVR on new requirement variations. The results show that TVR still maintains 97.13% in accuracy, demonstrating strong robustness. Additionally, to address the challenge of missing links, we applied TVR to traceability link recovery, successfully identifying 502 missing traceability links with an accuracy of 85.50%.
Though we focus on automotive software and more specifically safety-critical DTC requirements, many aspects of our approach can a priori be adapted to other contexts. For example, automotive systems must usually manage errors at run-time due to hardware failures or other causes. This is typically regulated by standards (Siegl et al., 2010; ISO/IEC 33001, 5 03; ISO/IEC 33002, 5 03; ISO/IEC 33003, 5 03; ISO/IEC 33004, 5 03), including traceability aspects.
@@
  • We propose TVR, a RAG-enhanced LLM approach that retrieves similar requirement pairs to be included in prompts, for validating traceability links between automotive stakeholder and system requirements.
  • We evaluate and demonstrate the robustness of TVR on unseen requirement variations, since it is common for requirements writing conventions to vary in industrial contexts.
  • We also apply TVR to recover traceability links and obtain high accuracy.
  • We comprehensively investigate 13 SOTA LLMs across five prompt engineering strategies, as well as the impact of different similarity measures and the number of examples on TVR performance.
@@ 2.1.2. DTC Stakeholder Requirements
Figure 1. Fictitious Variations of Stakeholder Requirements (Due to data privacy concerns, only fictitious examples are shown here. Italicized variables in the examples (e.g., MESSAGE_1, SIGNAL_1) are pseudonyms—real were employed in the experiments.)
@@ 2.1.3. DTC System Requirements
Figure 2. Example of Sanitized DTC System Requirement. Note: Real values were employed in the experiments.
@@ 2.2. Requirements Traceability
To address these two challenges, this paper proposes TVR, an approach for verifying requirements traceability for automotive systems. TVR leverages the exceptional natural language understanding capabilities of LLMs. Instead of relying on the general concept of traceability links (Guo et al., 2024), we ask LLMs to verify whether our system requirement covers the message or signal specified in the stakeholder requirement. We guide the LLM to focus specifically on the message or signal, rather than other parts of the requirement, to avoid confusion.
@@ 3. TVR
The overall workflow of TVR, illustrated in Figure 3, consists of two key components: a retriever and a generator. The retriever identifies the most similar examples from $D_{\text{train}}$ based on their similarity to $x_{\text{test}}$ and provides them as input to the generator. The generator then utilizes these retrieved examples to assess the validity of the traceability link between $stakeReq_{\text{test}}$ and $sysReq_{\text{test}}$.
@@ 3. RQ3:
    How robust is TVR for requirements traceability validation in the presence of unseen requirement variations?
    One of our main motivations in relying on LLMs is that we expect them to be more robust to such variations. Due to the numerous variations that typically present in stakeholder requirements, this RQ primarily evaluates the performance of TVR across unseen variations in stakeholder requirements. In practice, this is important as we expect to continuously encounter new variations.
@@ 4.2.1. LLMs
In this study, due to our industry partner’s data privacy concerns, we were unable to use LLMs such as the GPT series (OpenAI, 2023a, b). Instead, we were provided with 13 SOTA LLMs available through the Amazon Bedrock service, which include Claude, Llama, Mistral, and Titan (as detailed in Table 1). Extensive experiments were conducted using these LLMs.
Table 1. Evaluated Large Language Models.
Model Name Release Date Para- meters Max Tokens Supported Use Cases
Claude 3.5 Sonnet 24.06 - 200K RAG or search & retrieval over vast amounts of knowledge, product recommendations, forecasting, targeted marketing, code generation, quality control, parse text from images.
Claude 3 Sonnet 24.03 - 200K RAG or search & retrieval over vast amounts of knowledge, product recommendations, forecasting, targeted marketing, code generation, quality control, parse text from images.
Claude 2 23.07 - 100K Thoughtful dialogue, content creation, complex reasoning, creativity, and coding.
Claude Instant 23.08 - 100K Casual dialogue, text analysis, summarization, and document comprehension.
Claude 3 Haiku 24.03 - 200K Quick and accurate support in live interactions, translations, content moderation, optimize logistics, inventory management, extract knowledge from unstructured data.
Llama 3 8B 24.04 8B 8K Text summarization, text classification, sentiment analysis, and language translation.
Llama 3 70B 24.04 70B 8K Text summarization and accuracy, text classification and nuance, sentiment analysis and nuance reasoning, language modeling, dialogue systems, code generation, and following instructions.
Mistral 7B 24.03 7B 32K Text summarization, structuration, question answering, and code completion.
Mixtral 8x7B 24.03 46.7B 32K Text summarization, structuration, question answering, and code completion.
Mistral Large 2402 24.07 123B 128K Multilingual translation, text summarization, complex multilingual reasoning tasks, math and coding tasks including code generation.
Amazon Titan Text Premier 24.05 - 32K Chat, chain of thought reasoning, open-ended text generation, brainstorming, summarization, code generation, table creation, data formatting, paraphrasing, rewriting, extraction, and Q&A
Amazon Titan Text Express 23.11 - 8K Open-ended text generation, brainstorming, summarization, code generation, table creation, data formatting, paraphrasing, chain of thought, rewriting, extraction, Q&A, and chat
Amazon Titan Text Lite 24.09 - 4K Summarization and copywriting.
@@ 4.2.2. Prompt Engineering
  • Zero-shot Prompting (Brown et al., 2020). LLMs, pre-trained on vast amounts of data, can perform tasks in a zero-shot manner, meaning they require no prior task-specific examples. Zero-shot prompting directly instructs the model to complete a task without providing demonstrations or additional context.
  • Chain-of-Thought Prompting (CoT) (Wei et al., 2022). CoT enhances the complex reasoning capabilities of LLMs by incorporating intermediate reasoning steps. It has been widely used for tasks requiring arithmetic, commonsense, and symbolic reasoning, improving the model’s ability to handle multi-step problems.
@@
  • Retrieval-Augmented Generation (RAG) (Lewis et al., 2020). RAG enhances LLMs’ capability of solving knowledge-intensive tasks by integrating information retrieval with text generation. It retrieves relevant external knowledge sources to provide additional context, enabling models to generate more accurate and up-to-date responses without requiring full retraining (Gao et al., 2023b; Riedel et al., 2020). This makes RAG particularly effective for tasks involving fact-based and evolving knowledge.
Note that in our experiments, each prompting strategy builds upon the previous one. The CoT reasoning extends Zero-Shot, while few-shot prompting enhances CoT by incorporating examples in the prompt. The self-consistency strategy further refines few-shot prompting by applying a majority voting approach. RAG improves few-shot prompting by selecting more relevant examples.
@@ 4.3.1. Data Collection
To evaluate the performance of LLMs in automotive requirements traceability validation and recovery, we collected a dataset of DTC requirements from our industry partner. We began by gathering high-quality DTC system requirements, including the “Lost Communication” and “Implausible Data” DTC system requirements, which are among the most common and critical types of DTC requirements. Subsequently, we identified the stakeholder requirements linked to these system requirements, yielding a total of 1,320 stakeholder requirements.
@@ 4.3.2. Data Cleaning
To check the completeness of stakeholder requirements, we defined regular expressions to validate their structure  [^8] While we acknowledge that more advanced approaches for completeness checking exist, this is not the primary focus of our study.. Specifically, each atomic requirement must include a trigger condition, an input message, and a mitigation action. We filtered out both stakeholder and system requirements that did not meet the above two constraints. In practice, such requirements should be flagged as issues requiring correction and must be addressed before performing requirements traceability validation.
@@ 4.3.3. Data Annotation
Two authors independently labeled all 2,132 traceability links between stakeholder and system requirements, following the system engineers’ guidelines. Any discrepancies were discussed and resolved until consensus was achieved. When necessary, a system engineer was consulted to provide expert guidance and ensure the accuracy of the manual labeling. Ultimately, 1,913 links were confirmed as valid, while 219 were identified as invalid, indicating no traceability relationship should exist between those pairs.
4.3.4. Data Privacy
Since our dataset is obtained directly from our industry partner’s live production environment, we are therefore subject to strict confidentiality obligations: No details about the data may be disclosed under any circumstances. All labeling and experimentation must be conducted exclusively on company-issued laptops, and the data must never be copied, transferred, or stored outside these secure devices. Additionally, although we use the industry partner’s licensed LLM services for our work, we are expressly prohibited from processing the data on any third-party or public LLM platforms (e.g., ChatGPT) to eliminate any risk of unauthorized exposure.
@@ 4.4.1. Traceability Validation
To assess the performance of TVR in automotive requirements traceability validation, we employ standard metrics, including accuracy, precision, recall, and F1 score. (see Formula 1, 2, 3 and 4). $TP$ is the number of traceability links that are correctly identified as valid. $FN$ is the number of traceability links that are incorrectly identified as invalid. $FP$ is the number of traceability links that are incorrectly identified as valid. $TN$ is the number of traceability links that are correctly identified as invalid.
(1) ${Accuracy=\frac{TP+TN}{TP+TN+FP+FN}}$
(2) ${Precision=\frac{TP}{TP+FP}}$
(3) ${Recall=\frac{TP}{TP+FN}}$
(4) ${F1=\frac{2\times Precision\times Recall}{Precision+Recall}}$
As shown in Formula 1, accuracy is the proportion of correctly identified traceability links (both valid and invalid) out of all traceability links. Precision (Formula 2) is the proportion of correctly identified valid links ($TP$) out of all links identified as valid by the model ($TP$ + $FP$). It measures how reliable the model’s predictions of valid links are. Recall (Formula 3) is the proportion of correctly identified valid links ($TP$) out of all actual valid links in the ground truth ($TP$ + $FN$). It measures the model’s ability to find all the valid links. F1 score (Formula 4) is the harmonic mean of precision and recall. It balances the trade-off between precision and recall, providing a single measure of the model’s performance.